# Utility Boundary of Dataset Distillation: Scaling and Configuration-Coverage Laws

### v1

@@ 1 Introduction
Dataset distillation (DD) (wang2018dataset; sucholutsky2021soft), also known as dataset condensation (DC) (zhao2020dataset; wang2022cafe), seeks to synthesize a compact dataset that enables models to approach the accuracy of full-data training while greatly reducing storage and compute costs. Over the past few years, three main categories of matching-based methods have emerged. Gradient matching (GM) aligns gradients between real and synthetic data through bilevel optimization, extended by augmentation consistency (zhao2021dataset), diversity regularization (cazenavette2023generalizing), and reverse matching (ye2024distilled). Distribution matching (DM) matches feature statistics, from early MMD-based formulations (li2017mmd) to higher-order or quantile-based variants (wang2022cafe; zhang2024m3d; wei2024dataset). Trajectory matching (TM) aligns full optimization dynamics, introduced in MTT (cazenavette2022dataset) and later extended to self-supervised and detection tasks (lee2023self; qi2024fetch) (a comprehensive review of related work is provided in Appendix B).
Despite empirical advances, the theoretical foundation of DD remains fragmented. Existing analyses are confined to paradigm-specific assumptions: GM theory is largely restricted to first-order gradient matching  (zhao2021dataset; deng2022remember); DM relies on kernel- or moment-based statistics (li2017mmd), thereby neglecting optimization dynamics; and TM, while empirically strong (cazenavette2022dataset), lack rigorous convergence guarantees beyond heuristic approximations. These paradigm-specific limitations highlight the absence of a unified theoretical view to relate different DD approaches, making it unclear why all three categories can yield near full-data utility or how to explain recurring empirical patterns such as the saturation of accuracy gains with larger distilled sample sizes (cazenavette2022dataset). A further challenge is robustness to training-configuration shifts, e.g., changes in optimizer, architecture, or augmentation between distillation and downstream training. Since DD is expected to substitute for the full dataset in practice, distilled data must remain effective under such shifts. Yet current evaluations often restrict to fixed setups or mild parameter perturbations (nguyen2021dataset; zhao2023dataset), and reported results reveal instability: sensitivity to random seeds (wang2018dataset), reliance on augmentation (zhao2021dataset), and weak cross-architecture transfer (liu2022dataset). To this, we introduce the notion of a utility boundary: the relationship between the distilled sample size and the diversity of configurations within which the distilled dataset can still match the performance of full-data training.
@@ 2 Preliminaries and Problem Setup
Single training configuration. Most theories are established under a single fixed training configuration (e.g., optimizer, hyperparameters, augmentation, architecture) and show that particular matching strategies (gradient, distribution, or trajectory) improve accuracy in the single setting (zhao2020dataset; cazenavette2022dataset; zhao2023dataset). Formally, one training step under a (fixed) configuration can be written as:
@@ 3 A Unified configuration-Dynamics-Error Framework
with feasible set $\Gamma_{a}$ for parameters. The way we update parameters here generalizes classical stochastic approximation and adaptive methods (robbins1951stochastic; bottou2018optimization). We consider a target configuration family $\mathcal{A}\subseteq\mathcal{C}$ of training configurations that reflects the intended deployment setting (shalev2014understanding).
@@
This metric is inspired by the stability and uniform convergence analyses of algorithmic dynamics (hardt2016train; raginsky2017non). The coverage complexity of $\mathcal{A}$ at radius $r>0$ under $d_{\mathcal{A}}$ is $\mathcal{H}_{\mathrm{cov}}(\mathcal{A},r)=\log N(\mathcal{A},d_{\mathcal{A}},r),$ where $N(\mathcal{A},d_{\mathcal{A}},r)$ is the minimal number of $d_{\mathcal{A}}$-balls of radius $r$ needed to cover $\mathcal{A}$. Here $r$ is mathematically the covering radius, which determines the resolution of ecological distinctions: smaller $r$ resolves finer differences and thus increases $\mathcal{H}_{\mathrm{cov}}(\mathcal{A},r)$.
@@
This notion unifies classical discrepancy measures used in dataset distillation and domain adaptation (zhao2020dataset; cazenavette2022dataset; zhao2023dataset; ben2006analysis). Specializing to empirical real and empirical synthetic distributions gives $\Delta_{a}(\hat{\mu}_{\tau},\hat{\mu}_{s})$.
Stability. The final stage of our configuration–dynamics–error framework concerns how discrepancies in dynamics translate into generalization error. Throughout the paper, our analysis builds on stability- and information-theoretic approaches to generalization (bousquet2002stability; russo2016controlling; xu2017information), and decomposes the error into three components: (i) an optimization residual, determined by the optimization steps $T$; (ii) statistical fluctuations, arising from finite sample sizes $n,m,k$; and (iii) a matching term, governed by the alignment discrepancy $\Delta_{a}$. This unified form can be summarized as
@@ 4 Single–configuration Generalization Bound
On the feasible parameter domain $\Gamma_{a}$, we assume: (i) bounded per–sample update and preconditioner, i.e., $\|g_{a}(\theta;z)\|\leq B_{g}$, $\|P_{a}(\theta)\|\leq\kappa_{a}$; (ii) bounded loss $|\ell(\theta;z)|\leq B_{\ell}$ and $L_{R}$–Lipschitz test risk $\hat{R}$; (iii) contractive dynamics under a PL–type condition, with contraction rate $\rho_{a}\in(0,1)$ and constant $C_{2,a}=(1-\rho_{a})/L_{R}$. These assumptions are standard in stability/generalization analyses and PL-based convergence (see hardt2016train; bousquet2002stability; karimi2016linear).
@@
where $e_{g}=\mathcal{O}(1/\sqrt{k}+1/\sqrt{n})$ is the fluctuation from distillation and training samples (see, e.g., bartlett2002rademacher, for Rademacher-based rates), and $e_{\mathrm{te}}=\mathcal{O}(1/\sqrt{m})$ is the test concentration error (see, e.g., vershynin2018high, for standard sub-Gaussian bounds). If distilled dataset $\mathcal{D}_{s}$ generate depending on real dataset $\mathcal{D}_{\tau}$, then $e_{g}$ further incurs an information–theoretic penalty $\mathcal{O}(\sqrt{I(\mathcal{D}_{s};\mathcal{D}_{\tau})/k})$ (russo2016controlling) (see Appendix C for proof details).
@@ 7 Experiments
We verify our theoretical results on MNIST (lecun2002gradient), CIFAR-10/100 (krizhevsky2009learning), and ImageNette (5206848). We evaluate four canonical distillation families: Gradient Matching (DC (zhao2020dataset), DSA (zhao2021dataset)), Distribution Matching (DM (zhao2023dataset)), Trajectory Matching (MTT (cazenavette2022dataset)), and a recent diffusion-based method (MGD3 (chansantiago2025mgd3modeguideddatasetdistillation)). Each target configuration consists of an architecture (ConvNet, LeNet (lecun2002gradient), ResNet-18 (he2016deep), AlexNet (krizhevsky2012imagenet)), optimizer (SGD/Adam), and augmentation (DSA on/off). The distilled dataset is always generated in the source configuration ConvNet+SGD, using the official open-source implementations released by the corresponding papers to ensure consistency and reproducibility across methods. We vary the number of distilled samples $k\in\{2,4,6,8,12,18,28,51,100,200\}$, and report the generalization error $\Delta_{a}(\hat{\mu}_{\tau},\hat{\mu}_{s})=\big|\hat{R}(\theta_{T}^{(s,a)})-\hat{R}(\theta_{T}^{(\tau,a)})\big|,$ where $a$ indexes the target configuration. In practice, we approximate the coverage complexity $\mathcal{H}_{\mathrm{cov}}(\mathcal{A},r)$ by $\log M$, where $M$ is the number of sampled configurations.
@@ 8 Conclusion and Limitations
Ethics Statement
This work follows the ICLR Code of Ethics and the broader principles of responsible research. It does not involve human subjects, personal data, or other sensitive information that would require ethics approval. All datasets used are publicly available and properly licensed, and we cite their original sources. To support transparency and reproducibility, we release our implementation and describe our experimental settings in detail. The research was conducted without conflicts of interest or external sponsorships that could have influenced its design, execution, or reporting.
@@ B.1 Methodological Advances in Dataset Distillation
Gradient Matching (GM). Dataset distillation was first formulated via gradient matching, where synthetic data are optimized to align the training gradients of real data (zhao2020dataset). Differentiable Siamese augmentation improved stability and generalization across pipelines (zhao2021dataset). Later works explored diversity regularization (cazenavette2022dataset), representative matching (DREAM) (liu2023dreammatch), and feature regression (FRePo) (zhou2022dataset). Engineering efforts such as DC-BENCH (cui2022dc) provided standardized evaluation. Despite progress, GM often struggles with cross-architecture transfer.
Distribution Matching (DM). Distribution-based methods align feature or embedding distributions rather than raw gradients. Zhao and Bilen (zhao2023dataset) proposed matching distributions through MMD in intermediate feature space. CAFe introduced hierarchical alignment (wang2022cafe), while M3D (zhang2024m3d) and latent quantile matching (LQM) (wei2024dataset) further improved statistical robustness. DM methods emphasize scalability and robustness, with strong performance under augmentation shifts.
Trajectory Matching (TM). Trajectory matching aligns the full optimization dynamics rather than single-step gradients. The MTT framework (cazenavette2022dataset) established this principle, later improved by truncated backpropagation (kim2022dataset) and automated trajectory design (liu2024dataset). Extensions include self-supervised distillation (zhou2024self). TM captures optimization pathways faithfully but is computationally demanding.
Beyond GM/DM/TM. Generative priors constrain synthetic samples in pretrained generative latent spaces, e.g., GLaD (cazenavette2023generalizing) and diffusion-based methods (su2024d). Large-scale efforts like SRe2(yin2023squeeze) and TESLA (cui2023scaling) scale distillation to ImageNet-1K. Beyond images, condensation has been extended to graphs (liu2024graph; zheng2023rdm) and adversarially robust regimes (sun2024diversity). Label-centric approaches argue that soft labels often dominate performance gains (sucholutsky2021soft; chen2023data). These directions illustrate the versatility of dataset distillation across domains and modalities. We do not directly evaluate generative-based methods in our experiments, since our focus is on canonical optimization-driven procedures (GM/DM/TM). Nevertheless, our theoretical framework is general and applies equally to settings where synthetic data are produced by generative priors or large-scale latent models, as the alignment–risk perspective only depends on the induced empirical distributions rather than the mechanism of synthesis.
@@ B.2 Theoretical Explorations in Dataset Distillation
Kernel and NTK perspectives. Several works analyze distillation under kernel ridge regression or NTK approximations. KIP (nguyen2020dataset) and its infinite-width extension (nguyen2021dataset) provided conditions for exact recovery. Random feature approximation (RFAD) improved scalability (loo2022efficient). Recent works extend these guarantees to provable bounds (chen2024provable).
Generalization and stability. Classical results on uniform stability (hardt2016train), information-theoretic generalization (xu2017information; bu2020tightening), and Rademacher complexity (bartlett2002rademacher) have been adapted to study synthetic datasets. Convexified implicit gradient methods (wai2020accelerating) provide insights into bilevel optimization stability, directly relevant to GM and TM.
Scaling laws and soft labels. Empirical analyses consistently show saturation as the number of distilled samples (IPC) increases (cazenavette2022dataset; zhao2023dataset). Soft-label studies (sucholutsky2021soft; chen2023data) demonstrate Pareto frontiers relating IPC and generalization accuracy, suggesting the existence of an irreducible alignment floor.
configuration and transfer. Cross-architecture transferability remains a challenge: distilled datasets often fail under mismatched architectures or augmentations (liu2023dreammatch; cazenavette2023generalizing). Large-scale efforts like SRe2(yin2023squeeze) and TESLA (cui2023scaling), as well as generative priors (cazenavette2023generalizing; su2024d), can be interpreted as attempts to reduce alignment floors or expand coverage. Design space analyses (shao2024elucidating) and diversity–realism tradeoff studies (sun2024diversity) further quantify ecological robustness.
@@ C.1 Technical lemmas used in the proofs
Condition Eq. (C.1) is implied, e.g., by a uniform gradient bound $\sup_{\xi\in\Gamma_{a}}\|\nabla_{\theta}\ell(\xi;z)\|\leq L_{R}$ for all $z$ via the mean value theorem (e.g., bubeck2015convex; nesterov2013introductory). It also immediately implies that every risk functional $R_{\nu}(\theta)=\mathbb{E}_{\nu}\ell(\theta;z)$—for any probability measure $\nu$ on zz—is $L_{R}$–Lipschitz, since expectations preserve Lipschitz moduli (see Step 2 below). Thus, Assumption 4.1 plus Eq. (C.1) yields Lipschitzness uniformly over $\nu$, including $\nu=\hat{\nu}$.
@@
This is the standard contraction property of gradient descent under smoothness and strong convexity (see nesterov2013introductory; bubeck2015convex) or under the PL condition (see karimi2016linear). Applying it with $\mu=\hat{\mu}_{s}$ and $(\theta,\theta^{\prime})=(\theta_{t}^{(s)},\theta_{t}^{(\tau)})$ gives
@@
We use the (empirical) Rademacher complexity with the factor “2” built in (mohri2018foundations): for a sample $U=(z_{1},\dots,z_{m})$,
@@
Under this normalization, the standard uniform deviation bound (mohri2018foundations; see also bartlett2002rademacher; xu2016local) reads: for any $\varepsilon\in(0,1)$, with probability at least $1-\varepsilon$,
@@
(The $\widehat{\mathfrak{R}}_{U}$ term can be further upper bounded by $\mathfrak{R}_{m}(\mathcal{F}_{a})$ in expectation; we will retain $\mathfrak{R}_{m}(\cdot)$ notation below. A direct treatment of the vector norm via the vector-contraction inequality maurer2016vector, theorem 3 yields the same dependence on the scalarized class.)
@@
Replacing the empirical complexities by their expectations gives a sample-independent version with $\mathfrak{R}_{k}(\mathcal{F}_{a})+\mathfrak{R}_{n}(\mathcal{F}_{a})$. If one uses the “no-2” normalization for Rademacher complexity (mohri2018foundations), the bound incurs the standard extra factor $2$ in front of $\mathfrak{R}_{m}^{\mathrm{std}}(\mathcal{F}_{a})$. ∎
@@
Standard empirical process bound via symmetrization and concentration (mohri2018foundations; see also bartlett2002rademacher).
@@
Mutual-information (MI) tail inequality. The only difference of this proof compared with lemma C.4 is that we assumed istilled dataset $\mathcal{D}_{s}$ depends on $\mathcal{D}_{\tau}$. Thus, we apply a high-probability MI generalization bound (bu2020tightening; cf. xu2017information, steinke2020reasoning): for any $\varepsilon_{\mathrm{mi}}\in(0,1)$, with probability at least $1-\varepsilon_{\mathrm{mi}}$,
@@
Iterations $T$. Assume the PL-type decay $\rho_{a}^{T}\leq C_{\mathrm{opt}}/T^{\beta_{\mathrm{opt}}}$ (karimi2016linear). To enforce $L_{R}\rho_{a}^{T}\|\delta_{0}\|\leq\beta_{0}\epsilon_{0}$, it suffices that
@@
Distilled samples $k$. Under standard rates for Rademacher complexity (bartlett2002rademacher; xu2016local; mohri2018foundations), assume there exist $C_{g},C_{\mathrm{te}}$ such that $\mathfrak{R}_{k}(\mathcal{G}_{a})\leq C_{g}/\sqrt{k}$, $\mathfrak{R}_{n}(\mathcal{G}_{a})\leq C_{g}/\sqrt{n}$. Then

### v3

@@ 1 Introduction
Dataset distillation (DD) (Wang et al., 2018; Sucholutsky & Schonlau, 2021), also known as dataset condensation (DC) (Zhao et al., 2020; Wang et al., 2022), seeks to synthesize a compact dataset that enables models to approach the accuracy of full-data training while greatly reducing storage and compute costs. Over the past few years, three main categories of matching-based methods have emerged. Gradient matching (GM) aligns gradients between real and synthetic data through bilevel optimization, extended by augmentation consistency (Zhao & Bilen, 2021), diversity regularization (Cazenavette et al., 2023), and reverse matching (Ye et al., 2024). Distribution matching (DM) matches feature statistics, from early MMD-based formulations (Li et al., 2017) to higher-order or quantile-based variants (Wang et al., 2022; Zhang et al., 2024; Wei et al., 2024). Trajectory matching (TM) aligns full optimization dynamics, introduced in MTT (Cazenavette et al., 2022) and later extended to self-supervised and detection tasks (Lee et al., 2023; Qi et al., 2024) (a comprehensive review of related work is provided in Appendix B).
Despite empirical advances, the theoretical foundation of DD remains fragmented. Existing analyses are confined to paradigm-specific assumptions: GM theory is largely restricted to first-order gradient matching  (Zhao & Bilen, 2021; Deng & Russakovsky, 2022); DM relies on kernel- or moment-based statistics (Li et al., 2017), thereby neglecting optimization dynamics; and TM, while empirically strong (Cazenavette et al., 2022), lack rigorous convergence guarantees beyond heuristic approximations. These paradigm-specific limitations highlight the absence of a unified theoretical view to relate different DD approaches, making it unclear why all three categories can yield near full-data utility or how to explain recurring empirical patterns such as the saturation of accuracy gains with larger distilled sample sizes (Cazenavette et al., 2022). A further challenge is robustness to training-configuration shifts, e.g., changes in optimizer, architecture, or augmentation between distillation and downstream training. Since DD is expected to substitute for the full dataset in practice, distilled data must remain effective under such shifts. Yet current evaluations often restrict to fixed setups or mild parameter perturbations (Nguyen et al., 2021; Zhao & Bilen, 2023), and reported results reveal instability: sensitivity to random seeds (Wang et al., 2018), reliance on augmentation (Zhao & Bilen, 2021), and weak cross-architecture transfer (Liu et al., 2022). To this, we introduce the notion of a utility boundary: the relationship between the distilled sample size and the diversity of configurations within which the distilled dataset can still match the performance of full-data training.
@@ 2 Preliminaries and Problem Setup
Single training configuration. Most theories are established under a single fixed training configuration (e.g., optimizer, hyperparameters, augmentation, architecture) and show that particular matching strategies (gradient, distribution, or trajectory) improve accuracy in the single setting (Zhao et al., 2020; Cazenavette et al., 2022; Zhao & Bilen, 2023). Formally, one training step under a (fixed) configuration can be written as:
@@ 3 A Unified configuration-Dynamics-Error Framework
with feasible set $\Gamma_{a}$ for parameters. The way we update parameters here generalizes classical stochastic approximation and adaptive methods (Robbins & Monro, 1951; Bottou et al., 2018). We consider a target configuration family $\mathcal{A}\subseteq\mathcal{C}$ of training configurations that reflects the intended deployment setting (Shalev-Shwartz & Ben-David, 2014).
@@
This metric is inspired by the stability and uniform convergence analyses of algorithmic dynamics (Hardt et al., 2016; Raginsky et al., 2017). The coverage complexity of $\mathcal{A}$ at radius $r>0$ under $d_{\mathcal{A}}$ is $\mathcal{H}_{\mathrm{cov}}(\mathcal{A},r)=\log N(\mathcal{A},d_{\mathcal{A}},r),$ where $N(\mathcal{A},d_{\mathcal{A}},r)$ is the minimal number of $d_{\mathcal{A}}$-balls of radius $r$ needed to cover $\mathcal{A}$. Here $r$ is mathematically the covering radius, which determines the resolution of ecological distinctions: smaller $r$ resolves finer differences and thus increases $\mathcal{H}_{\mathrm{cov}}(\mathcal{A},r)$.
@@
This notion unifies classical discrepancy measures used in dataset distillation and domain adaptation (Zhao et al., 2020; Cazenavette et al., 2022; Zhao & Bilen, 2023; Ben-David et al., 2006). Specializing to empirical real and empirical synthetic distributions gives $\Delta_{a}(\hat{\mu}_{\tau},\hat{\mu}_{s})$.
Stability. The final stage of our configuration–dynamics–error framework concerns how discrepancies in dynamics translate into generalization error. Throughout the paper, our analysis builds on stability- and information-theoretic approaches to generalization (Bousquet & Elisseeff, 2002; Russo & Zou, 2016; Xu & Raginsky, 2017), and decomposes the error into three components: (i) an optimization residual, determined by the optimization steps $T$; (ii) statistical fluctuations, arising from finite sample sizes $n,m,k$; and (iii) a matching term, governed by the alignment discrepancy $\Delta_{a}$. This unified form can be summarized as
@@ 4 Single–configuration Generalization Bound
On the feasible parameter domain $\Gamma_{a}$, we assume: (i) bounded per–sample update and preconditioner, i.e., $\|g_{a}(\theta;z)\|\leq B_{g}$, $\|P_{a}(\theta)\|\leq\kappa_{a}$; (ii) bounded loss $|\ell(\theta;z)|\leq B_{\ell}$ and $L_{R}$–Lipschitz test risk $\hat{R}$; (iii) contractive dynamics under a PL–type condition, with contraction rate $\rho_{a}\in(0,1)$ and constant $C_{2,a}=(1-\rho_{a})/L_{R}$. These assumptions are standard in stability/generalization analyses and PL-based convergence (see Hardt et al. (2016); Bousquet & Elisseeff (2002); Karimi et al. (2016)).
@@
where $e_{g}=\mathcal{O}(1/\sqrt{k}+1/\sqrt{n})$ is the fluctuation from distillation and training samples (see, e.g., Bartlett & Mendelson (2002), for Rademacher-based rates), and $e_{\mathrm{te}}=\mathcal{O}(1/\sqrt{m})$ is the test concentration error (see, e.g., Vershynin (2018), for standard sub-Gaussian bounds). If distilled dataset $\mathcal{D}_{s}$ generate depending on real dataset $\mathcal{D}_{\tau}$, then $e_{g}$ further incurs an information–theoretic penalty $\mathcal{O}(\sqrt{I(\mathcal{D}_{s};\mathcal{D}_{\tau})/k})$ (Russo & Zou, 2016) (see Appendix C for proof details).
@@ 7 Experiments
We verify our theoretical results on MNIST (LeCun et al., 2002), CIFAR-10/100 (Krizhevsky et al., 2009), and ImageNette (Deng et al., 2009). We evaluate four canonical distillation families: Gradient Matching (DC (Zhao et al., 2020), DSA (Zhao & Bilen, 2021)), Distribution Matching (DM (Zhao & Bilen, 2023)), Trajectory Matching (MTT (Cazenavette et al., 2022)), and a recent diffusion-based method (MGD3 (Chan-Santiago et al., 2025)). Each target configuration consists of an architecture (ConvNet, LeNet (LeCun et al., 2002), ResNet-18 (He et al., 2016), AlexNet (Krizhevsky et al., 2012)), optimizer (SGD/Adam), and augmentation (DSA on/off). The distilled dataset is always generated in the source configuration ConvNet+SGD, using the official open-source implementations released by the corresponding papers to ensure consistency and reproducibility across methods. We vary the number of distilled samples $k\in\{2,4,6,8,12,18,28,51,100,200\}$, and report the generalization error $\Delta_{a}(\hat{\mu}_{\tau},\hat{\mu}_{s})=\big|\hat{R}(\theta_{T}^{(s,a)})-\hat{R}(\theta_{T}^{(\tau,a)})\big|,$ where $a$ indexes the target configuration. In practice, we approximate the coverage complexity $\mathcal{H}_{\mathrm{cov}}(\mathcal{A},r)$ by $\log M$, where $M$ is the number of sampled configurations.
@@ B.1 Methodological Advances in Dataset Distillation
Gradient Matching (GM). Dataset distillation was first formulated via gradient matching, where synthetic data are optimized to align the training gradients of real data (Zhao et al., 2020). Differentiable Siamese augmentation improved stability and generalization across pipelines (Zhao & Bilen, 2021). Later works explored diversity regularization (Cazenavette et al., 2022), representative matching (DREAM) (Liu et al., 2023), and feature regression (FRePo) (Zhou et al., 2022). Engineering efforts such as DC-BENCH (Cui et al., 2022) provided standardized evaluation. Despite progress, GM often struggles with cross-architecture transfer.
Distribution Matching (DM). Distribution-based methods align feature or embedding distributions rather than raw gradients. Zhao and Bilen (Zhao & Bilen, 2023) proposed matching distributions through MMD in intermediate feature space. CAFe introduced hierarchical alignment (Wang et al., 2022), while M3D (Zhang et al., 2024) and latent quantile matching (LQM) (Wei et al., 2024) further improved statistical robustness. DM methods emphasize scalability and robustness, with strong performance under augmentation shifts.
Trajectory Matching (TM). Trajectory matching aligns the full optimization dynamics rather than single-step gradients. The MTT framework (Cazenavette et al., 2022) established this principle, later improved by truncated backpropagation (Kim et al., 2022) and automated trajectory design (Liu et al., 2024a). Extensions include self-supervised distillation (Zhou et al., 2024). TM captures optimization pathways faithfully but is computationally demanding.
Beyond GM/DM/TM. Generative priors constrain synthetic samples in pretrained generative latent spaces, e.g., GLaD (Cazenavette et al., 2023) and diffusion-based methods (Su et al., 2024). Large-scale efforts like SRe2(Yin et al., 2023) and TESLA (Cui et al., 2023) scale distillation to ImageNet-1K. Beyond images, condensation has been extended to graphs (Liu et al., 2024b; Zheng & Li, 2023) and adversarially robust regimes (Sun et al., 2024). Label-centric approaches argue that soft labels often dominate performance gains (Sucholutsky & Schonlau, 2021; Chen et al., 2023). These directions illustrate the versatility of dataset distillation across domains and modalities. We do not directly evaluate generative-based methods in our experiments, since our focus is on canonical optimization-driven procedures (GM/DM/TM). Nevertheless, our theoretical framework is general and applies equally to settings where synthetic data are produced by generative priors or large-scale latent models, as the alignment–risk perspective only depends on the induced empirical distributions rather than the mechanism of synthesis.
@@ B.2 Theoretical Explorations in Dataset Distillation
Kernel and NTK perspectives. Several works analyze distillation under kernel ridge regression or NTK approximations. KIP (Nguyen et al., 2020) and its infinite-width extension (Nguyen et al., 2021) provided conditions for exact recovery. Random feature approximation (RFAD) improved scalability (Loo et al., 2022). Recent works extend these guarantees to provable bounds (Chen et al., 2024).
Generalization and stability. Classical results on uniform stability (Hardt et al., 2016), information-theoretic generalization (Xu & Raginsky, 2017; Bu et al., 2020), and Rademacher complexity (Bartlett & Mendelson, 2002) have been adapted to study synthetic datasets. Convexified implicit gradient methods (Wai et al., 2020) provide insights into bilevel optimization stability, directly relevant to GM and TM.
Scaling laws and soft labels. Empirical analyses consistently show saturation as the number of distilled samples (IPC) increases (Cazenavette et al., 2022; Zhao & Bilen, 2023). Soft-label studies (Sucholutsky & Schonlau, 2021; Chen et al., 2023) demonstrate Pareto frontiers relating IPC and generalization accuracy, suggesting the existence of an irreducible alignment floor.
configuration and transfer. Cross-architecture transferability remains a challenge: distilled datasets often fail under mismatched architectures or augmentations (Liu et al., 2023; Cazenavette et al., 2023). Large-scale efforts like SRe2(Yin et al., 2023) and TESLA (Cui et al., 2023), as well as generative priors (Cazenavette et al., 2023; Su et al., 2024), can be interpreted as attempts to reduce alignment floors or expand coverage. Design space analyses (Shao et al., 2024) and diversity–realism tradeoff studies (Sun et al., 2024) further quantify ecological robustness.
@@ C.1 Technical lemmas used in the proofs
Condition Eq. (C.1) is implied, e.g., by a uniform gradient bound $\sup_{\xi\in\Gamma_{a}}\|\nabla_{\theta}\ell(\xi;z)\|\leq L_{R}$ for all $z$ via the mean value theorem (e.g., Bubeck et al. 2015, Prop. B.10; Nesterov 2013, Chap. 2). It also immediately implies that every risk functional $R_{\nu}(\theta)=\mathbb{E}_{\nu}\ell(\theta;z)$—for any probability measure $\nu$ on zz—is $L_{R}$–Lipschitz, since expectations preserve Lipschitz moduli (see Step 2 below). Thus, Assumption 4.1 plus Eq. (C.1) yields Lipschitzness uniformly over $\nu$, including $\nu=\hat{\nu}$.
@@
This is the standard contraction property of gradient descent under smoothness and strong convexity (see Nesterov 2013, theorem 2.1.15; Bubeck et al. 2015, Prop. 3.5) or under the PL condition (see Karimi et al. 2016, theorem 1). Applying it with $\mu=\hat{\mu}_{s}$ and $(\theta,\theta^{\prime})=(\theta_{t}^{(s)},\theta_{t}^{(\tau)})$ gives
@@
We use the (empirical) Rademacher complexity with the factor “2” built in (Mohri et al. 2018, Def. 3.1): for a sample $U=(z_{1},\dots,z_{m})$,
@@
Under this normalization, the standard uniform deviation bound (Mohri et al. 2018, theorem 3.3; see also Bartlett & Mendelson 2002; Xu et al. 2016) reads: for any $\varepsilon\in(0,1)$, with probability at least $1-\varepsilon$,
@@
(The $\widehat{\mathfrak{R}}_{U}$ term can be further upper bounded by $\mathfrak{R}_{m}(\mathcal{F}_{a})$ in expectation; we will retain $\mathfrak{R}_{m}(\cdot)$ notation below. A direct treatment of the vector norm via the vector-contraction inequality Maurer, 2016, theorem 3 yields the same dependence on the scalarized class.)
@@
Replacing the empirical complexities by their expectations gives a sample-independent version with $\mathfrak{R}_{k}(\mathcal{F}_{a})+\mathfrak{R}_{n}(\mathcal{F}_{a})$. If one uses the “no-2” normalization for Rademacher complexity (Mohri et al., 2018), the bound incurs the standard extra factor $2$ in front of $\mathfrak{R}_{m}^{\mathrm{std}}(\mathcal{F}_{a})$. ∎
@@
Standard empirical process bound via symmetrization and concentration (Mohri et al. 2018, theorem 3.3; see also Bartlett & Mendelson 2002).
@@
Mutual-information (MI) tail inequality. The only difference of this proof compared with lemma C.4 is that we assumed istilled dataset $\mathcal{D}_{s}$ depends on $\mathcal{D}_{\tau}$. Thus, we apply a high-probability MI generalization bound (Bu et al. 2020, theorem 7; cf. Xu & Raginsky 2017, theorem 1, Steinke & Zakynthinou 2020, theorem 1): for any $\varepsilon_{\mathrm{mi}}\in(0,1)$, with probability at least $1-\varepsilon_{\mathrm{mi}}$,
@@
Iterations $T$. Assume the PL-type decay $\rho_{a}^{T}\leq C_{\mathrm{opt}}/T^{\beta_{\mathrm{opt}}}$ (Karimi et al. 2016, Theorem 1). To enforce $L_{R}\rho_{a}^{T}\|\delta_{0}\|\leq\beta_{0}\epsilon_{0}$, it suffices that
@@
Distilled samples $k$. Under standard rates for Rademacher complexity (Bartlett & Mendelson 2002; Xu et al. 2016; Mohri et al. 2018), assume there exist $C_{g},C_{\mathrm{te}}$ such that $\mathfrak{R}_{k}(\mathcal{G}_{a})\leq C_{g}/\sqrt{k}$, $\mathfrak{R}_{n}(\mathcal{G}_{a})\leq C_{g}/\sqrt{n}$. Then